### v2

@@ Abstract
Mamba, a recent selective structured state space model, excels in long sequence modeling, which is vital in the large model era. Long sequence modeling poses significant challenges, including capturing long-range dependencies within the data and handling the computational demands caused by their extensive length. Mamba addresses these challenges by overcoming the local perception limitations of convolutional neural networks and the quadratic computational complexity of Transformers. Given its advantages over these mainstream foundation architectures, Mamba exhibits great potential to be a visual foundation architecture. Since January 2024, Mamba has been actively applied to diverse computer vision tasks, yielding numerous contributions. To help keep pace with the rapid advancements, this paper reviews visual Mamba approaches, analyzing over 200 papers. This paper begins by delineating the formulation of the original Mamba model. Subsequently, it delves into representative backbone networks, and applications categorized using different modalities, including image, video, point cloud, and multi-modal. Particularly, we identify scanning techniques as critical for adapting Mamba to vision tasks, and decouple these scanning techniques to clarify their functionality and enhance their flexibility across various applications. Finally, we discuss the challenges and future directions, providing insights into new outlooks in this fast evolving area. A comprehensive list of visual Mamba models reviewed in this work is available at https://github.com/Ruixxxx/Awesome-Vision-Mamba-Models.
Keywords:
Mamba State Space Model Computer Vision Application
1 Introduction
@@
Artificial intelligence technologies, especially deep learning, have revolutionized numerous application fields. In the field of computer vision (CV), convolutional neural networks (CNNs) utilize local receptive fields and shared weights to process visual data, capitalizing on inductive biases such as locality and spatial invariance (Krizhevsky et al, 2012; Simonyan and Zisserman, 2015; He et al, 2016). Despite their efficient linear computational complexity with respect to image resolution and proficiency in modeling local patterns, CNNs have restricted receptive fields. This restriction limits their capability to capture larger spatial contexts, which is essential for comprehensively understanding scenes or complex spatial relations in tasks that demand global information. In recent years, Vision Transformers (ViTs) (Dosovitskiy et al, 2021), which utilize a self-attention mechanism (Vaswani et al, 2017) to process sequences of image patches, have demonstrated remarkable modeling capabilities across various visual tasks (Liu et al, 2021). Self-attention enables ViTs to capture long-range dependencies within images, providing a significant advantage over traditional CNNs that rely on local receptive fields. This capability allows ViTs to exhibit robust performance on varied datasets and scale effectively to large model sizes. However, the self-attention mechanism involves a quadratic computational cost to the number of patches, which limits the scalability of ViTs. The CV domain has long been dominated by CNNs and ViTs, each with their respective strengths and inherent limitations. To overcome their limitations, researchers have devoted significant effort to improving these models. Recently, structured state space models (Gu et al, 2022b, 2023) have garnered considerable attention due to their computational efficiency and principled capability in modeling long-range dependencies (Gu et al, 2020).
@@
Figure 1: The statistics of Mamba-based papers released to date on vision tasks, spanning different modalities including Image, Video, Point Cloud, and Multi-Modal
@@
The state space model is a concept that is widely adopted in various disciplines. Its core idea is connecting the input and output sequences using a latent state. It takes different forms in different disciplines, such as Markov decision process in reinforcement learning (Hafner et al, 2020), dynamic causal modeling in computational neuroscience (Friston et al, 2003) and Kalman filters in controls (Kalman, 1960). Recently, the state space model (SSM) has been introduced to deep learning for sequence modeling and its parameters or mappings are learned by gradient descent (Gu et al, 2021). SSM is essentially a type of sequence transformation and can be incorporated into deep neural networks. It conceptually unifies the strengths of former paradigms for sequence model design, including continuous-time models (CTMs), recurrent neural networks (RNNs), and CNNs. Thereinto, although CNNs are primarily used for processing spatial data and are not sequential, they can be adapted for sequence modeling by leveraging their abilities to encode local contexts and facilitate parallelizable computations. However, SSMs have not been widely used in practice due to their extensive computational and memory requirements coming from the state representation. This situation has changed with the advent of the structured SSM (S4), which addresses these limitations through reparameterizing the state matrices (Gu et al, 2022b). Since then, a series of SSM transformations and neural network architectures incorporating SSM layers emerge (Gu et al, 2022b, a; Gupta et al, 2022; Orvieto et al, 2023; Smith et al, 2023; Gu et al, 2023; Hasani et al, 2023; Fu et al, 2023). However, SSMs’ constant sequence transitions restrict their context-based reasoning ability, which is important for the efficacy of models like Transformer (Vaswani et al, 2017). In Mamba (Gu and Dao, 2023), the authors propose to address this by integrating a selection mechanism into the SSM, thus enabling the SSM to selectively propagate or forget information along the sequence or scan path based on the current token. Furthermore, to efficiently compute these selective SSMs, the authors develop a hardware-aware algorithm. Subsequently, the authors integrate these selective SSMs into a simplified neural network architecture, termed Mamba. With the modeling power akin to Transformers and linear scalability in sequence length, Mamba becomes a promising foundation architecture for sequence modeling, which is crucial in the large model era.
@@
Due to the growing adoption of techniques from sequence modeling or natural language processing (NLP) into CV, there is a rapid application of Mamba (Gu and Dao, 2023) to CV tasks (Zhang et al, 2024a; Liu et al, 2024f; Heidari et al, 2024; Zou et al, 2024b). VMamba (Liu et al, 2024g) is an early representative visual Mamba model. It unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. Another visual Mamba model Vim (Zhu et al, 2024a) leverages position embeddings to incorporate spatial information, inspired by ViT (Dosovitskiy et al, 2021). It also uses bi-directional SSM for handling the non-causal image sequences. Similarly, several other notable studies (Li et al, 2024d; Yang et al, 2024a; Patro and Agneeswaran, 2024b; Huang et al, 2024e; Pei et al, 2024) delve into the exploration of visual backbone networks, consistently achieving competitive performance across classification, detection, and segmentation tasks. To highlight the balance between efficiency and effectiveness in visual Mamba models, Fig. 5 offers a graphical representation contrasting their performance with computational complexity. In addition to these efforts, Mamba has been applied across a diverse range of vision modalities and their respective applications, encompassing image processing, video analysis, point cloud processing, and multi-modal scenarios. Endowed with the aforementioned modeling capability and linear scalability, Mamba stands out as a promising foundation architecture for CV tasks. The growing interest among researchers in applying Mamba to various vision tasks is reflected in the increasing number of studies dedicated to this exploration, as plotted in Fig. 1.
@@
In the rapidly evolving field of CV, Mamba (Gu and Dao, 2023) has emerged as a significant advancement. Keeping pace with the latest research is critical for the community. Therefore, this paper aims to provide a comprehensive review of the applications of Mamba in visual tasks, shedding light on both its foundational elements and diverse applications across various modalities. Our contributions are summarized as follows:
@@
    Formulation of Mamba (Section 2): We provide an introductory overview of the operational principles of the Mamba (Gu and Dao, 2023) and highlight its distinctions from traditional state space models.
@@
    Applications (Section 4): We categorize applications of Mamba by different modalities, such as image, video, point cloud, and multi-modal data. Each category is explored in depth to highlight how the Mamba framework adapts to and benefits individual modalities.
@@
2 Formulation of Mamba
Mamba (Gu and Dao, 2023) is a recent sequence model aiming at improving the context-based reasoning ability of state space model (SSM) by simply making its parameters to be functions of the input. The SSM here especially refers to the sequence transformation used in the structured state space sequence model (S4) (Gu et al, 2022b), which can be incorporated into deep neural networks. Mamba simplifies the commonly used SSM block and forms a simplified SSM architecture. In the following, we elaborate on the core concepts of Mamba.
2.1 State Space Model
The SSM transformation in S4 (Gu et al, 2022b) originates from the classical state space model, which maps a 1D input signal $x(t)\in\mathbb{R}$ to a 1D output signal $y(t)\in\mathbb{R}$ through an N-D latent state $h(t)\in\mathbb{R}^{N}$. This transformation is mathematically formulated as linear ordinary differential equations (ODEs):
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t),$
@@ 2.1 State Space Model
To deal with the discrete input sequence $\boldsymbol{x}=(x_{0},x_{1},...)\in\mathbb{R}^{L}$, various discretization rules can be used to discretize the parameters in Eq. (1) using a step size $\Delta$, which can be seen as the resolution of the continuous input $x(t)$. Following previous work (Tustin, 1947), S4 (Gu et al, 2022b) discretizes these parameters using the bilinear method. Following (Gupta et al, 2022), Mamba (Gu and Dao, 2023) employs the zero-order hold (ZOH) assumption [^1] ZOH assumes that the sample value of $x$ remains constant over each sampling interval $\Delta$. For more details on ZOH, refer to https://en.wikipedia.org/wiki/Zero-order_hold to solve the ODEs. Specifically, this results in the conversion of the continuous parameters $\mathbf{A},\mathbf{B}$ into their discrete counterparts $\overline{\mathbf{A}},\overline{\mathbf{B}}$ as follows:
$\overline{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),$ (2)
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.$
@@
$h_{t}$ $=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},$ (3)
$y_{t}$ $=\mathbf{C}h_{t}.$
@@
For efficient parallelizable training, this recursive process can also be reformulated and computed as a convolution (Gu et al, 2022b):
$\overline{\mathbf{K}}$ $=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),$ (4)
$\boldsymbol{y}$ $=\boldsymbol{x}*\overline{\mathbf{K}},$
@@
Figure 2: Mamba block: a simplified block that integrates the H3 and MLP blocks
2.2 Selective SSM
As seen, the parameters in SSM indicated by either Eq. (1), Eq. (3) or Eq. (4) remain invariant with respect to the input or temporal dynamics. Mamba (Gu and Dao, 2023) identifies this linear time-invariant (LTI) property as a fundamental limitation of SSM when it comes to context-based reasoning. To address this issue, Mamba incorporates a selection mechanism. The selection mechanism is implemented by simply configuring the parameters of SSM as functions of the input, thus achieving input-dependent interactions along the sequence. Specifically, parameters $\mathbf{B},\mathbf{C},\Delta$ are dependent on the input sequence $\boldsymbol{x}$:
$\mathbf{B}$ $=s_{B}(\boldsymbol{x}),$ (5)
$\mathbf{C}$ $=s_{C}(\boldsymbol{x}),$
$\Delta$ $=\tau_{\Delta}(Parameter+s_{\Delta}(\boldsymbol{x}).$
@@ 2.2 Selective SSM
The resulting selective SSM cannot be computed as either RNN or convolution. Mamba (Gu and Dao, 2023) employs a hardware-aware algorithm to efficiently compute the selective SSM. The hardware-aware algorithm leverages three classical techniques: parallel scan, kernel fusion, and recomputation.
2.3 Mamba

### v3

@@ Abstract.
Mamba, a recent selective structured state space model, excels in long sequence modeling, which is vital in the large model era. Long sequence modeling poses significant challenges, including capturing long-range dependencies within the data and handling the computational demands caused by their extensive length. Mamba addresses these challenges by overcoming the local perception limitations of convolutional neural networks and the quadratic computational complexity of Transformers. Given its advantages over these mainstream foundation architectures, Mamba exhibits great potential to be a visual foundation architecture. Since January 2024, Mamba has been actively applied to diverse computer vision tasks, yielding numerous contributions. To help keep pace with the rapid advancements, this paper reviews visual Mamba approaches, analyzing nearly 300 papers. This paper begins by delineating the formulation of the original Mamba model. Subsequently, it delves into representative backbone networks, and applications categorized using different modalities, including image, video, point cloud, and multi-modal data. Particularly, we identify scanning techniques as critical for adapting Mamba to vision tasks, and decouple these scanning techniques to clarify their functionality and enhance their flexibility across various applications. Finally, we discuss the challenges and future directions, providing insights into new outlooks in this fast evolving area. A comprehensive list of visual Mamba models reviewed in this work is available at https://github.com/Ruixxxx/Awesome-Vision-Mamba-Models.
Mamba, State Space Model, Computer Vision, Application
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Artificial intelligence technologies, especially deep learning, have revolutionized numerous application fields. In the field of computer vision (CV), convolutional neural networks (CNNs) utilize local receptive fields and shared weights to process visual data, capitalizing on inductive biases such as locality and spatial invariance (Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; He et al., 2016). Despite their efficient linear computational complexity with respect to image resolution and proficiency in modeling local patterns, CNNs have restricted receptive fields. This restriction limits their capability to capture larger spatial contexts, which is essential for comprehensively understanding scenes or complex spatial relations in tasks that demand global information. In recent years, Vision Transformers (ViTs) (Dosovitskiy et al., 2021), which utilize a self-attention mechanism (Vaswani et al., 2017) to process sequences of image patches, have demonstrated remarkable modeling capabilities across various visual tasks (Liu et al., 2021). Self-attention enables ViTs to capture long-range dependencies within images, providing a significant advantage over traditional CNNs that rely on local receptive fields. This capability allows ViTs to exhibit robust performance on varied datasets and scale effectively to large model sizes. However, the self-attention mechanism involves a quadratic computational cost to the number of patches, which limits the scalability of ViTs. The CV domain has long been dominated by CNNs and ViTs, each with their respective strengths and inherent limitations. To overcome their limitations, researchers have devoted significant effort to improving these models. Recently, structured state space models (Gu et al., 2022a, 2023) have garnered considerable attention due to their computational efficiency and principled capability in modeling long-range dependencies (Gu et al., 2020).
@@
Figure 1. The statistics of Mamba-based papers released to date on vision tasks, spanning different modalities including Image, Video, Point Cloud, and Multi-Modal.
[…]
@@
The state space model is a concept that is widely adopted in various disciplines. Its core idea is connecting the input and output sequences using a latent state. It takes different forms in different disciplines, such as Markov decision process in reinforcement learning (Hafner et al., 2020), dynamic causal modeling in computational neuroscience (Friston et al., 2003) and Kalman filters in controls (Kalman, 1960). Recently, the state space model (SSM) has been introduced to deep learning for sequence modeling and its parameters or mappings are learned by gradient descent (Gu et al., 2021). SSM is essentially a type of sequence transformation and can be incorporated into deep neural networks. It conceptually unifies the strengths of former paradigms for sequence model design, including continuous-time models (CTMs), recurrent neural networks (RNNs), and CNNs. Among these, although CNNs are primarily used for processing spatial data and are not sequential, they can be adapted for sequence modeling by leveraging their abilities to encode local contexts and facilitate parallelizable computations. However, SSMs have not been widely used in practice due to their extensive computational and memory requirements coming from the state representation. This situation has changed with the advent of the structured SSM (S4), which addresses these limitations through reparameterizing the state matrices (Gu et al., 2022a). Since then, a series of SSM transformations and neural network architectures incorporating SSM layers emerge (Gu et al., 2022a, b; Gupta et al., 2022; Orvieto et al., 2023; Smith et al., 2023; Gu et al., 2023; Hasani et al., 2023; Fu et al., 2023). However, SSMs’ constant sequence transitions restrict their context-based reasoning ability, which is important for the efficacy of models like Transformer (Vaswani et al., 2017). In Mamba-1 (Gu and Dao, 2024), the authors propose to address this by integrating a selection mechanism into the SSM, thus enabling the SSM to selectively propagate or forget information along the sequence or scan path based on the current token. Furthermore, to efficiently compute these selective SSMs, the authors develop a hardware-aware algorithm. In Mamba-2 (Dao and Gu, 2024a), the authors establish connections between the SSMs and variants of attention, and refine the selective SSM into a state space duality (SSD) algorithm to further improve computational efficiency. With the modeling power akin to Transformers and linear scalability in sequence length, Mamba becomes a promising foundation architecture for sequence modeling, which is crucial in the large model era.
@@
Due to the growing adoption of techniques from sequence modeling or natural language processing (NLP) into CV, there is a rapid application of Mamba to CV tasks (Zhang et al., 2024i; Liu et al., 2024i; Heidari et al., 2024; Zou et al., 2024a). VMamba (Liu et al., 2024e), an early representative visual Mamba model, unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. Another visual Mamba model Vim (Zhu et al., 2024c) leverages position embeddings to incorporate spatial information, inspired by ViT (Dosovitskiy et al., 2021). It also uses bi-directional SSM for handling the non-causal image sequences. Similarly, several other notable studies (Li et al., 2024h; Yang et al., 2024a; Patro and Agneeswaran, 2024b; Huang et al., 2024d; Pei et al., 2024) delve into the exploration of visual backbone networks, consistently achieving competitive performance across classification, detection, and segmentation tasks. To highlight the balance between efficiency and effectiveness in visual Mamba models, Fig. 5 offers a graphical representation contrasting their performance with computational complexity. In addition to these efforts, Mamba has been applied across a diverse range of vision modalities and their respective applications, encompassing image processing, video analysis, point cloud processing, and multi-modal scenarios. Endowed with the aforementioned modeling capability and linear scalability, Mamba stands out as a promising foundation architecture for CV tasks. The growing interest among researchers in applying Mamba to various vision tasks is reflected in the increasing number of studies dedicated to this exploration, as plotted in Fig. 1.
@@
In the rapidly evolving field of CV, Mamba has emerged as a significant advancement. Keeping pace with the latest research is critical for the community. Therefore, this paper aims to provide a comprehensive review of the applications of Mamba in vision tasks, shedding light on both its foundational elements and diverse applications across various modalities. Our contributions are summarized as follows:
@@
    Formulation of Mamba (Section 2): We provide an introductory overview of the operational principles of the Mamba (Gu and Dao, 2024; Dao and Gu, 2024a) and highlight its distinctions from traditional state space models.
@@
    Applications (Section 4): We categorize applications of Mamba by different modalities, such as image, video, point cloud, and multi-modal data. Each category is explored in depth to highlight how the Mamba architecture adapts to and benefits individual modalities.
@@
2. Formulation of Mamba
Mamba (Gu and Dao, 2024) is a recent sequence model aiming at improving the context-based reasoning ability of state space model (SSM) by simply making its parameters to be functions of the input. The SSM here especially refers to the sequence transformation used in the structured state space sequence model (S4) (Gu et al., 2022a), which can be incorporated into deep neural networks. Mamba-2 (Dao and Gu, 2024b) further provides a theoretical framework that connects structured SSMs and variants of attention, thereby enabling the transfer of algorithmic and system optimizations originally developed for Transformers to SSMs. In the following, we elaborate on the core concepts of Mamba-1 and Mamba-2.
2.1. State Space Model
The SSM transformation in S4 (Gu et al., 2022a) originates from the classical state space model, which maps a 1D input signal $x(t)\in\mathbb{R}$ to a 1D output signal $y(t)\in\mathbb{R}$ through an N-D latent state $h(t)\in\mathbb{R}^{N}$. This transformation is mathematically formulated as linear ordinary differential equations (ODEs):
(1) $h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$
@@ 2.1. State Space Model
To deal with the discrete input sequence $\boldsymbol{x}=(x_{0},x_{1},...)\in\mathbb{R}^{L}$, various discretization rules can be used to discretize the parameters in Eq. (1) using a step size $\Delta$, which can be seen as the resolution of the continuous input $x(t)$. Following previous work (Tustin, 1947), S4 (Gu et al., 2022a) discretizes these parameters using the bilinear method. Following (Gupta et al., 2022), Mamba (Gu and Dao, 2024) employs the zero-order hold (ZOH) assumption [^1] ZOH assumes that the sample value of $x$ remains constant over each sampling interval $\Delta$. For more details on ZOH, refer to https://en.wikipedia.org/wiki/Zero-order_hold to solve the ODEs. Specifically, this results in the conversion of the continuous parameters $\mathbf{A},\mathbf{B}$ into their discrete counterparts $\overline{\mathbf{A}},\overline{\mathbf{B}}$ as follows:
(2) $\overline{\mathbf{A}}=\exp(\Delta\mathbf{A}),\quad\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.$
@@
(3) $h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t}.$
@@
For efficient parallelizable training, this recursive process can be reformulated and computed as a convolution (Gu et al., 2022a):
(4) $\overline{\mathbf{K}}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),\quad\boldsymbol{y}=\boldsymbol{x}*\overline{\mathbf{K}},$
@@
2.2. Mamba-1: Selective SSM
As seen, the parameters in SSM indicated by either Eq. (1), Eq. (3) or Eq. (4) remain invariant with respect to the input or temporal dynamics. Mamba (Gu and Dao, 2024) identifies this linear time-invariant (LTI) property as a fundamental limitation of SSM when it comes to context-based reasoning. To address this issue, Mamba incorporates a selection mechanism. The selection mechanism is implemented by simply configuring the parameters of SSM as functions of the input, thus achieving input-dependent interactions along the sequence. Specifically, parameters $\mathbf{B},\mathbf{C},\Delta$ are dependent on the input sequence $\boldsymbol{x}$:
(5) $\mathbf{B}=s_{B}(\boldsymbol{x}),\quad\mathbf{C}=s_{C}(\boldsymbol{x}),\quad\Delta=\tau_{\Delta}(Parameter+s_{\Delta}(\boldsymbol{x})),$
@@ 2.2. Mamba-1: Selective SSM
The resulting selective SSM cannot be computed as either RNN or convolution. Mamba (Gu and Dao, 2024) employs a hardware-aware algorithm to efficiently compute the selective SSM. The hardware-aware algorithm leverages three classical techniques: parallel scan, kernel fusion, and recomputation.
2.3. Mamba-2: State Space Duality
To leverage the collective advancements made by the community on Transformers, Mamba-2 (Dao and Gu, 2024a) first reformulates the SSM transformation through induction as:
(6) $y_{t}=\sum_{s=0}^{t}\mathbf{C}_{t}^{\top}\mathbf{A}_{t:s}^{\times}\mathbf{B}_{s}x_{s},$
where $\mathbf{A}_{t:s}^{\times}=\mathbf{A}_{t}\times\cdots\times\mathbf{A}_{s+1},s<t$. The matrix transformation form of SSM can then be derived as:
(7) $\boldsymbol{y}=\text{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\boldsymbol{x})=\mathbf{M}\boldsymbol{x},$
where $\mathbf{M}_{ji}=\mathbf{C}_{j}^{\top}\mathbf{A}_{j}\cdots\mathbf{A}_{i+1}\mathbf{B}_{i}$. Consequently, SSM also has a quadratic algorithm besides the linear algorithm.
Mamba-2 considers a special case of structured SSM when $\mathbf{A}_{j}$ is a scalar. Then the matrix can be rearranged as:
(8) $\mathbf{M}_{ji}=\mathbf{A}_{j:i}\cdot(\mathbf{C}_{j}^{\top}\mathbf{B}_{i}).$
This can be vectorized into:
(9) $\mathbf{M}=\mathbf{L}\circ(\mathbf{C}\mathbf{B}^{\top}),$
which is consistent with the original definition of masked kernel attention. To sum up, SSM with a scalar-identity structure in the $\mathbf{A}$ matrix, and structured masked attention with a 1-semiseparable structure in the $\mathbf{L}$ mask are dual to each other, exhibiting identical linear and quadratic forms.
Mamba-2 introduces an SSD algorithm based on block decompositions of $\mathbf{M}$, leveraging both the linear SSM recurrence and its quadratic dual form to improve computational and memory efficiency while utilizing matrix multiplication units on modern hardware. This SSD algorithm achieves a $2$-8×8$\times$8 × speedup over the hardware-aware Mamba-1 implementation.
2.4. Mamba Architecture